OpenStudy (anonymous):

is the integral of du/(3u+3) = (1/3)(du/(u+3))??? plsss help

OpenStudy (john_es):

You mean, \[\int \frac{du}{3u+3}=\frac{1}{3}\int \frac{du}{u+1}\]

OpenStudy (anonymous):

yes.

Miracrown (miracrown):